# Equivariant concordance of periodic 2-knots in $S^{4}$

### v1

@@ Abstract.
We show that the equivariant concordance group of smooth 2-knots in $S^{4}$ invariant under a $\mathbb{Z}/d\mathbb{Z}$ action, where the action is given by rotation about an unknotted sphere intersecting the 2-knot in two points, is isomorphic to $\mathbb{Z}/2\mathbb{Z}$ for all $d\geq 2$. This is in contrast to the non-equivariant case, in which all 2-knots are slice. We construct a new invariant for these 2-knots, which we call periodic, and show that it fully classifies them up to equivariant concordance. The invariant depends on an extension of the Arf invariant for null-homologous classical knots in arbitrary 3-manifolds.
@@ Introduction
Concordance of surfaces in various contexts has been an object of interest to topologists as early as the 1960s. In 1965, Kervaire [Ker65] showed that all 2-spheres in $S^{4}$ are slice, meaning that they bound smoothly embedded 3-balls into $B^{5}$. In 2015, Sunukjian [Sun15] generalized this to the statement that the concordance class of any oriented surface in a simply-connected 4-manifold is determined entirely by its genus and homology class. Our work uses many of the same techniques as these earlier papers, and extends the theory of concordance of 2-spheres to a $\mathbb{Z}/d\mathbb{Z}$-equivariant setting.
Our goal is to examine equivariant concordance of smooth 2-spheres that are symmetric with respect to an ambient rotation $\tau$ acting on $S^{4}$ fixing an unknotted $S^{2}$ of fixed finite order $d$, where the rotation restricted to the symmetric 2-sphere fixes exactly two points. In analogy with the classical dimension, we call such 2-knots $d$-periodic. We say that these spheres are equivariantly slice if they bound 3-balls into $B^{5}$ that are themselves equivariant with respect the extension of $\tau$ to $B^{5}$ that fixes an unknotted properly embedded 3-ball. Note that any such slice ball must intersect the fixed set of $\tau$ extended to $B^{5}$ in a single unknotted arc by Morgan’s resolution of the Smith Conjecture [Mor84]. There is also a notion of periodic connected sum, analogous to the connected sum of strongly invertible knots defined by Sakuma [Sak86], which turns the set of smooth concordance classes of periodic 2-knots up to equivariant isotopy into a group, denoted $\mathcal{C}^{2}_{d}$. This paper introduces a new invariant $\alpha$ of $d$-periodic 2-knots which completely determines their equivariant concordance class, simultaneously computing of all of these groups.
@@
There exists a $\mathbb{Z}/2\mathbb{Z}$-valued, additive concordance invariant $\alpha$ for $d$-periodic 2-knots $S$ such that:
@@
    if $\alpha(S)=0$, then $S$ is equivariantly slice,
@@
    if $\alpha(S)=1$, then $S$ is equivariantly concordant to the $d$-twist spun trefoil, and
@@
    every $d$-periodic 2-knot satisfies exactly one of these conditions.
@@
Thus the smooth equivariant concordance group $\mathcal{C}^{2}_{d}$ is isomorphic to $\mathbb{Z}/2\mathbb{Z}$ for all $d$.
@@
We prove our main theorem as follows. We first construct equivariant slice balls by constructing equivariant Seifert 3-manifolds, pushing these into $B^{5}$, and then equivariantly surgering them to 3-balls. The surgery approach fails under certain conditions, which we then show to be due to an actual equivariant concordance obstruction. This obstruction relies on a new extension of the Arf invariant for knots to null-homologous knots in arbitrary 3-manifolds, detailed in Section 3. Realizing this invariant as an equivariant sliceness obstruction requires us to also build equivariant Seifert 4-manifolds for periodic 3-manifolds in ambient 5-space with careful control of the fixed set.
It is worth emphasizing that the concordance group, $\mathbb{Z}/2\mathbb{Z}$, does not depend on the period $d$ of the rotation. In short, this is because the equivariant concordance of periodic spheres depends only on the isotopy class of the quotient sphere relative to its intersection with the axis of rotation. We thus obstruct this type of relative concordance between these quotient spheres, and include criteria for when this concordance can be lifted to an equivariant concordance on the cover. Thus our proof really shows that the concordance group of annuli properly embedded in $S^{1}\times B^{3}$ is $\mathbb{Z}/2\mathbb{Z}$, explaining the lack of dependence on $d$. Another consequence of our proof is that the natural map from $\mathcal{C}^{2}_{kd}$ to $\mathcal{C}^{2}_{d}$ for $k\in\mathbb{N}$ induced by the inclusion of $kd$-periodic 2-knots into the set of $d$-periodic 2-knots is always an isomorphism, which further reflects insignificance of the period.
@@
This result fits into a body of work both on equivariant 2-knots and on surface concordance, in various categories. Ruberman [Rub84] showed that 2-knots which are equivariant with respect to free involutions on $S^{4}$ are unique up to equivariant concordance. Gabbard [Gab24] has also worked on 2-spheres that are also equivariant with respect to a rotation involution on $S^{4}$, but which are embedded in such a way that the rotation induces an orientation-reversing involution on the 2-sphere fixing a circle.
@@
Related also is the work of McCooey [McC07] on topological concordance of $\mathbb{Z}/p\mathbb{Z}\times\mathbb{Z}/p\mathbb{Z}$ actions on $S^{4}$ for primes $p$, where he also found a $\mathbb{Z}/2\mathbb{Z}$-valued action concordance invariant. McCooey’s definition of concordance of actions implies concordance of their fixed sets, and so in proving his theorem, he shows that there are at most two topological concordance classes of pairs of 2-knots in $S^{4}$ intersecting in two points, often called Montesinos twins, each of which are the fixed set of a different commuting $\mathbb{Z}/p\mathbb{Z}$ action on $S^{4}$. His proof uses surgery theory, and his obstruction comes from the generator of the $\mathbb{Z}/2\mathbb{Z}$ part of $L_{4}(\mathbb{Z}\times\mathbb{Z})\cong\mathbb{Z}\oplus\mathbb{Z}/2\mathbb{Z}$.
@@
Our work is substantially different from that of McCooey, although we conjecture that our obstruction to smooth equivariant sliceness, restricted to the topological category and the correct class of periodic 2-knots, is equivalent to McCooey’s. For one, McCooey’s result only has implications for equivariant topological concordance within the class of 2-knots which are not only $\mathbb{Z}/p\mathbb{Z}$-equivariant, but are also themselves the fixed sets of $\mathbb{Z}/p\mathbb{Z}$ actions on $S^{4}$ for $p$ prime. As mentioned before, his proof is fully surgery-theoretic, and he was unable to find an example of a $\mathbb{Z}/p\mathbb{Z}\times\mathbb{Z}/p\mathbb{Z}$ action on $S^{4}$ for which his obstruction vanished. In contrast, our theorem applies to all smooth periodic 2-knots of any period. The proof is completely constructive, and we give explicit criteria for the vanishing of our obstruction and identify it as a direct extension of the Arf invariant of a knot, as well as examples where it does not vanish. Although the Arf invariant is typically a topological invariant, we work in the smooth category for the remainder of this paper.
@@
There is also a family of literature on surface concordance in non-simply connected 4-manifolds, in which there are several known obstructions. In 2018 Gabai gave a proof of the 4-dimensional lightbulb theorem [Gab20], which gives conditions under which a sphere with a framed, geometrically dual sphere can be unknotted. Soon after, Miller proved a version for sphere concordance in the presence of framed duals [Mil21]. Both of these proofs require that the Freedman-Quinn invariant of the given 2-knot vanish in the case that the ambient 4-manifold has 2-torsion in its fundamental group. The connection to the Freedman-Quinn invariant was formalized by Schneiderman and Teichner [ST22] after a counterexample to the lightbulb theorem in the case of ambient 2-torsion was constructed by Schwartz [Sch19]. In the case that a sphere in a non-simply connected 4-manifold has an unframed dual, Klug and Miller [KM21] extended work of Stong [Sto93] and Freedman and Quinn [FQ90] to show that the Kervaire-Milnor invariant of the sphere can be used to obstruct concordance. Although we do not use these obstructions, we mention them in order to give a sense of scope of surface concordance invariants.
@@
  • In Section 1, we establish some preliminaries of our setup and outline some necessary results regarding equivariant Morse theory with vertical boundary. This requires the language of Morse theory for manifolds with boundary, as described by Borodzik-Némethi-Ranicki in [BNR16], as well as work of Borodzik-Powell in [BP16] and Schwartz in [Sch80].
  • In Section 2, we prove the existence of equivariant Seifert 3-manifolds for equivariant 2-knots while controlling the fixed set by adapting a technique of Yanagawa in [Yan69].
  • In Section 3, we take an interlude to describe an extension of the Arf invariant to null-homologous knots in 3-manifolds, inspired by a definition of Robertello in [Rob65].
  • In Section 4, we use the extension in Section 3 to define the $\alpha$ invariant for $d$-periodic 2-knots together with their Seifert solids, and then give a procedure to produce equivariant slice balls via surgery on these Seifert solids under the assumption that it vanishes. We also show that all knots for which $\alpha=1$ are also mutually equivariantly concordant.
  • In Section 5, we use the invariants defined in the previous sections to demonstrate that $\alpha$ is an invariant of $d$-periodic 2-knots up to equivariant isotopy, and that it obstructs equivariant sliceness. The proof involves the construction of certain equivariant Seifert 4-manifolds for 3-manifolds with $\mathbb{Z}/d\mathbb{Z}$ actions fixing a circle.
@@ Acknowledgments
The author would like to thank Maggie Miller, Imogen Montague, Mark Powell, Evan Scott, Aru Mukherjea, and Seungwon Kim for their generous help and advice. Special thanks to Malcolm Gabbard for suggesting this question, and to Mark Powell for his helpful comments on an earlier draft of this paper.
1. Preliminaries
1.1. Main definitions and conventions
In this paper, we work entirely in the smooth category. It is important for consistency to fix the following basic convention: when we say that $S$ is a 2-knot in $S^{4}$, we mean that $S$ is the image of a smooth embedding $\psi\colon S^{2}\to S^{4}$, considered up to smooth isotopy of $\psi$.
@@ 1.1. Main definitions and conventions
An action $\tau$ of $\mathbb{Z}/d\mathbb{Z}$ on the standard $m$-ball $B^{m}$ is linear if it is conjugate to the action of an $m$-dimensional orthogonal representation of $\mathbb{Z}/dZ$ on $B^{m}$. An action of $\mathbb{Z}/d\mathbb{Z}$ on $S^{m-1}$ is linear if it is conjugate to the action of such a representation on $B^{m}$ restricted to the boundary.
The actions we consider in this paper are the the linear rotation action $\rho$ of $\mathbb{Z}/d\mathbb{Z}$ on $S^{2}$ fixing two points, and the linear rotation action $\tau$ of $\mathbb{Z}/d\mathbb{Z}$ on $B^{5}$ fixing an unknotted 3-ball such that its restriction to the boundary $S^{4}$ fixes an unknotted 2-sphere. Restricting to linear actions simplifies the argument, although we hope to extend to nonlinear actions in future work. For convenience, we give a general definition for embeddings which are equivariant with respect to this type of action in all dimensions.
@@
Let $M$ be a smooth $n$-manifold and let $\psi\colon M\to S^{m}$ be a smooth embedding. We say that $\psi$ is $d$-periodic if there exists an order $d$ linear rotation $\tau$ on $S^{m}$ fixing an unknotted smooth copy of $S^{m-2}\subset S^{m}$ such that $\tau$ fixes the image of $\psi$ setwise and $\psi(M)$ intersects $\operatorname{Fix}(\tau)$ transversely.
In the case that $\psi(M)\cap\operatorname{Fix}(\tau)$ is a single, smoothly embedded copy of $S^{n-2}$, or a single properly embedded copy of $D^{n-2}$ in the case that $M$ has boundary, we say that $\psi$ is simply $d$-periodic.
Some abuses of this notation: where it is not explicitly relevant, we will often drop the prefix $d$- and describe embeddings just as periodic or simply periodic. Furthermore, we will also often describe the image of periodic embeddings as periodic; i.e., when we say that a 2-knot $S\subset S^{4}$ is periodic, we mean that there is an embedding $\psi\colon S^{2}\to S^{4}$ whose image is $S$, and that $\psi$ is periodic for some period $d$.
Following this definition, we see that a 2-knot $S$ in $S^{4}$ is $d$-periodic if it satisfies
$\tau\circ\psi=\psi\circ\rho,$
with $\operatorname{Fix}(\tau)$ an unknotted 2-sphere. In this case, since the restriction of $\tau$ to $S$ is orientation preserving, we see by the Lefschetz fixed point theorem that $S\cap\operatorname{Fix}(\tau)$ consists of exactly two points. Such spheres arise naturally in common constructions of 2-knots: all $d$-twist spun knots can be constructed to have this symmetry about their spin axis.
@@
All periodic 2-knots are simply periodic as noted above, but the same is not true of surface knots in $S^{4}$. For instance, the unknotted torus in $S^{4}$ admits two different order 2 periodic symmetries: one in which one of the $S^{1}$ factors of the torus forms a meridian for $\operatorname{Fix}(\tau)$, and another that restricts to the hyperelliptic involution. The former action is free, while the latter fixes four points.
Note that in principle, the intersection of a simply periodic manifold $M$ with the ambient fixed set can be knotted in $M$ (rel. boundary, if applicable). This will be of particular importance to us.
To aid in the exposition of the paper, we make some parametrization conventions. We think of our periodic 2-knot as being contained in the boundary of an upper half-ball $U\subset B^{5}$. We will frequently parametrize $U$ by the coordinates:
$U=\{(x,y,z,w,t)\mid x^{2}+y^{2}+z^{2}+w^{2}+t^{2}\leq C,t\geq 0\}$
for some $C>>1$. When we are doing constructions explicitly in $U\cap S^{4}$, we will refer to coordinates on $U$ simply as $(x,y,z,w)$. We also implicitly choose the standard metric on $U$, scaled appropriately to our choice of coordinates, which we use in our discussion of Morse theory. Submanifolds inside $U$ inherit this metric, and we will frequently differentiate between the gradient of a Morse function on the ambient space and the gradient of its restriction to the submanifold.
Since the action $\tau$ we consider on $B^{5}$ is the linear $\mathbb{Z}/d\mathbb{Z}$ action by rotation about a properly embedded 3-ball with unknotted boundary, we define our parameterization such that in coordinates, $\tau$ fixes the $(0,0,z,w,t)$-hyperplane in $U$.
We will also frequently need to pass between equivariant manifolds and their quotients by $\tau$. Throughout this paper, we denote the quotients of manifolds with overbars; e.g. $Y\subset B^{5}$ becomes $\overline{Y}\subset\overline{B^{5}}$ under $\tau$. (Because $\tau$ is always a linear cyclic action by rotation, the quotient of $B^{5}$ by $\tau$ is always diffeomorphic to $B^{5}$, and the quotient of $S^{4}$ by $\tau$ is always diffeomorphic to $S^{4}$.) To recover the original manifold, we may take the $d$-fold branched cover over the image of the fixed set under the covering map. For this reason, we refer to the image of $\operatorname{Fix}(\tau)$ under the covering map as the branch set and denote it $\operatorname{Branch}(\tau)$.
Finally, there are several points in this paper in which we need to refer to subsets of a manifold with a given Morse function on it. For this we use the following standard convention: if $f$ is a Morse function on a manifold $M$, then:
$M_{c}:=f^{-1}(c)$
for $c\in\mathbb{R}$. We also declare:
$M_{*c}:=f^{-1}(\{a\in\mathbb{R}\mid a*c\})$
where * is an element of ${<,,>,}\{$$\leq$$\geq\}$.
1.2. Equivariant concordance group structure
With these conventions settled, we may formally define the equivariant concordance group. This requires a notion of equivariant connected sum, which in turn requires some extra data. Equivariant connected sum for classical knots was first defined by Sakuma in [Sak86], and we follow the spirit of this construction in our own.

### v2

@@ Abstract.
We show that the smooth equivariant concordance group of 2-knots in $S^{4}$ invariant under a linear $\mathbb{Z}/d\mathbb{Z}$ action is isomorphic to $\mathbb{Z}/2\mathbb{Z}$ for all $d\geq 2$. This is in contrast to the non-equivariant case, in which all 2-knots are slice. We construct a new invariant for these 2-knots, which we call periodic, and show that it fully classifies them up to equivariant concordance. The invariant depends on a variation of the Arf invariant for null-homologous classical knots in arbitrary 3-manifolds with respect to a chosen spin structure. Our proof also shows an identical classification for certain annuli in $S^{1}\times B^{3}$ up to concordance rel. boundary.
@@ Introduction
Concordance of surfaces in various contexts has been an object of interest to topologists as early as the 1960s. In 1965, Kervaire [KER65] showed that all 2-spheres in $S^{4}$ are slice, meaning that they bound smoothly embedded 3-balls into $B^{5}$. In 2015, Sunukjian [SUN15] generalized this to the statement that the concordance class of any oriented surface in a simply-connected 4-manifold is determined entirely by its genus and homology class. Our work uses similar techniques as these earlier papers, and extends the theory of concordance of 2-spheres to a $\mathbb{Z}/d\mathbb{Z}$-equivariant setting. In doing so, we end up proving a classification theorem for certain annuli in $S^{1}\times B^{3}$ up to concordance which restricts to a product on the boundary.
Our goal is to examine equivariant concordance of smooth 2-spheres that are symmetric with respect to an ambient order $d$ rotation $\rho$ acting on $S^{4}$ fixing an unknotted $S^{2}$, where the rotation restricted to the symmetric 2-sphere fixes exactly two points. In analogy with the classical dimension, we call such 2-knots $d$-periodic. We say that these spheres are equivariantly slice if they bound 3-balls into $B^{5}$ that are themselves equivariant with respect the extension of $\rho$ to $B^{5}$ that fixes an unknotted properly embedded 3-ball. There is also a notion of periodic connected sum analogous to the connected sum of strongly invertible knots which turns the set of smooth equivariant concordance classes of periodic 2-knots into a group, denoted $\mathcal{C}^{d}_{2}$. This paper introduces a new invariant $\alpha_{d}$ of $d$-periodic 2-knots which completely determines their equivariant concordance class, simultaneously computing of all of these groups.
@@
For each positive integer $d$, there exists a $\mathbb{Z}/2\mathbb{Z}$-valued, additive concordance invariant $\alpha_{d}$ for $d$-periodic 2-knots $\tilde{S}$ such that:
@@
    if $\alpha_{d}(\tilde{S})=0$, then $\tilde{S}$ is $\mathbb{Z}/d\mathbb{Z}$-equivariantly slice,
@@
    if $\alpha_{d}(\tilde{S})=1$, then $\tilde{S}$ is $\mathbb{Z}/d\mathbb{Z}$-equivariantly concordant to the $d$-twist spun trefoil, and
@@
    every $d$-periodic 2-knot satisfies exactly one of the above two conditions.
@@
Thus the smooth equivariant concordance group $\mathcal{C}^{d}_{2}$ is isomorphic to $\mathbb{Z}/2\mathbb{Z}$ for all $d$.
We prove our main theorem by first proving an identical classification for annuli in $S^{1}\times B^{3}$ which are properly embedded such that both boundary components are copies of $S^{1}\times\text{pt}$ inside of $\partial(S^{1}\times B^{3})=S^{1}\times S^{2}$, which we call Montesinos annuli. We say that such an annulus is slice if it is concordant rel. boundary to a boundary parallel annulus in $S^{1}\times B^{3}$.
Theorem 5.2.
There exists a $\mathbb{Z}/2\mathbb{Z}$-valued, additive relative concordance invariant $\alpha$ for Montesinos annuli $N$ in $S^{1}\times B^{3}$ such that:
  1. (1)
    if $\alpha(N)=0$, then $N$ is slice,
  2. (2)
    if $\alpha(N)=1$, then $N$ is concordant to the one-twist spun trefoil minus its spin axis, and
  3. (3)
    every Montesinos annulus in $S^{1}\times B^{3}$ satisfies exactly one of the above two conditions.
Thus the smooth concordance group of Montesinos annuli $\mathcal{C}_{2}^{\circ}$ is isomorphic to $\mathbb{Z}/2\mathbb{Z}$.
@@
The argument for these theorems go as follows. Consider first a Montesinos annulus $N$. We can associate to $N$ a “corresponding” 2-sphere $S$ in $S^{4}$ which intersect a “removed” 2-sphere $R$ in two points, such that removing a neighborhood of $R$ would turn $S$ into $N$ inside $S^{1}\times B^{3}$. We construct a Seifert 3-manifold for this corresponding sphere which intersect the removed sphere in a single arc, interpreting this as a sort of relative Seifert manifold for the annulus. We then push this Seifert manifold into $B^{5}$, where it intersects a “removed” 3-ball $Q$ in a single arc, and then attempt to ambiently surger it to a 3-ball which intersect the removed 3-ball in a single unknotted arc. This ambient surgery approach fails under certain conditions, which we show depends on an extension of the Arf invariant for knots to null-homologous knots in arbitrary 3-manifolds, detailed in Section 2. Realizing this invariant as an annulus sliceness obstruction requires us to also build Seifert 4-manifolds for certain 3-manifolds in ambient 5-space with careful control of their intersection with an unknotted 3-sphere under specific Morse-theoretic conditions.
Once establishing this relative concordance group, we show that by lifting the ambient surgeries to the $d$-fold branched cover of $B^{5}$ over the removed 3-ball, we obtain a $\mathbb{Z}/d\mathbb{Z}$-equivariant slice ball for the branched cover of the sphere $S$ representing $N$ over its two intersection points with the removed sphere. In this way, we interpret the removed sphere as the unknotted fixed set of a linear rotation on $S^{4}$, and the annulus concordance as the quotient of the periodic concordance by the action $\rho$. This gives some intuition as to why the isomorphism type of these concordance groups, $\mathbb{Z}/2\mathbb{Z}$, does not depend on the period $d$ of the rotation. That said, the following question remains open:
Question 0.1.
Is the map from $\mathcal{C}^{kd}_{2}$ to $\mathcal{C}^{d}_{2}$ for $k\in\mathbb{N}$ induced by the inclusion of $kd$-periodic 2-knots into the set of $d$-periodic 2-knots is always an isomorphism? That is, does there exist a $kd$-periodic 2-knot which is not $\mathbb{Z}/kd\mathbb{Z}$-equivariantly slice, but is $\mathbb{Z}/d\mathbb{Z}$-equivariantly slice?
@@
This result fits into a body of work both on equivariant 2-knots and on surface concordance, in various categories. Ruberman [RUB84] showed that 2-knots which are equivariant with respect to free involutions on $S^{4}$ are unique up to equivariant concordance. Gabbard [GAB24] has also studied equivariant isotopy of 2-spheres that are also equivariant with respect to a rotation involution on $S^{4}$, but which are embedded in such a way that the rotation induces an orientation-reversing involution on the 2-sphere fixing a circle. Hellsten [HEL26] also has recent work on the classification of surfaces with boundary in simply-connected 4-manifolds up to concordance.
@@
Related also is the work of McCooey [MCC07] on topological concordance of $\mathbb{Z}/p\mathbb{Z}\times\mathbb{Z}/p\mathbb{Z}$ actions on $S^{4}$ for $p$ prime, where he also found a $\mathbb{Z}/2\mathbb{Z}$-valued action concordance invariant. McCooey’s definition of concordance of actions implies concordance of their fixed sets, and so in proving his theorem, he shows that there are at most two topological concordance classes of pairs of 2-knots in $S^{4}$ intersecting in two points, often called Montesinos twins, which are each the fixed set of a different commuting $\mathbb{Z}/p\mathbb{Z}$ action on $S^{4}$. His proof is surgery-theoretic, and the potential obstruction he identifies comes from the generator of the $\mathbb{Z}/2\mathbb{Z}$ part of $L_{4}(\mathbb{Z}[\mathbb{Z}\times\mathbb{Z}])\cong\mathbb{Z}\oplus\mathbb{Z}/2\mathbb{Z}$.
@@
Our work is substantially different from that of McCooey, reflecting the differing goals of his paper and ours, although we conjecture that our obstruction to smooth equivariant sliceness, restricted to the topological category and the correct class of periodic 2-knots, is equivalent to McCooey’s. Notably, McCooey’s result only has implications for equivariant topological concordance within the class of 2-knots which are not only $\mathbb{Z}/p\mathbb{Z}$-equivariant, but are also themselves the fixed sets of $\mathbb{Z}/p\mathbb{Z}$ actions on $S^{4}$ for $p$ prime. As mentioned before, his proof is fully surgery-theoretic, and it remains an open problem to find an example of a $\mathbb{Z}/p\mathbb{Z}\times\mathbb{Z}/p\mathbb{Z}$ action on $S^{4}$ for which his potential obstruction does not vanish. In contrast, our theorem applies to all smooth periodic 2-knots of any period. The proof is completely constructive, and we give explicit criteria for the vanishing of our obstruction and identify it as a direct extension of the Arf invariant of a knot, as well as examples where it does not vanish (although this does not immediately answer McCooey’s question on the existence of nonlinear $\mathbb{Z}/p\mathbb{Z}\times\mathbb{Z}/p\mathbb{Z}$ actions on $S^{4}$). Although the Arf invariant is typically a topological invariant, we work in the smooth category for the remainder of this paper.
@@
There is also a family of literature on surface concordance in non-simply connected 4-manifolds, in which there are several known obstructions. In 2018 Gabai gave a proof of the 4-dimensional lightbulb theorem [GAB20], which gives conditions under which a sphere with a framed, geometrically dual sphere can be unknotted. This result requires that the Freedman-Quinn invariant of the given sphere vanish in the case that the ambient 4-manifold has 2-torsion in its fundamental group. The connection to the Freedman-Quinn invariant was formalized by Schneiderman and Teichner [ST22] after a counterexample to the lightbulb theorem in the case of ambient 2-torsion was constructed by Schwartz [SCH19]. A similar result in the topological category had been proved by Freedman and Quinn [FQ90] showing that the Freedman-Quinn invariant obstructs concordance of 2-spheres, albeit with some errors which were corrected by Stong [STO93], and following Gabai’s work, Miller [MIL21] gave a constructive proof. In the case that a sphere in a non-simply connected 4-manifold has an unframed dual, Klug and Miller [KM21] extended work of Stong [STO93] and Freedman and Quinn [FQ90] to show that the Kervaire-Milnor invariant of the sphere can be used to obstruct concordance. Although we do not use these obstructions, we mention them in order to give a sense of scope of surface concordance invariants.
@@
In Section 1, we establish some preliminaries of our setup for both Montesinos annuli in $S^{1}\times B^{3}$ and periodic 2-knots. In Section 2, we describe an extension of the classical Arf invariant to null-homologous knots in 3-manifolds, adapting work of Klug [KLU21]. In Section 3, we outline some necessary results regarding embedded Morse theory with vertical boundary in the complement of a codimension two submanifold. This requires the language of Morse theory for manifolds with boundary, as described by Borodzik-Némethi-Ranicki in [BNR16], as well as work of Borodzik-Powell in [BP16], and inspired by work of Schwartz in [SCH80] in the equivariant setting. In Section 4, we establish the existence of Seifert 3-manifolds for 2-knots while controlling the intersection of these manifolds with an unknotted 2-sphere intersecting our knot in two points. We then construct Seifert 4-manifolds for 3-manifolds in $S^{5}$ while controlling the intersection of these manifolds with an unknotted 3-sphere, given certain Morse theoretic conditions. These both adapt a technique of Yanagawa in [YAN69]. In Section 5, we use the extension in Section 2 to define the $\alpha$ invariant for Montesinos annuli in $S^{1}\times B^{3}$, and then give a procedure to “slice” these annuli via surgery on the Seifert 3-manifolds constructed in Section 4 under the assumption that $\alpha$ vanishes. We also show that all Montesinos annuli for which $\alpha=1$ are also mutually relatively concordant. In Section 6, we use the results of the previous sections to define the invariant $\alpha_{d}$ for $d$-periodic 2-knots up to equivariant isotopy, and then lift all previous results to the equivariant setting.
@@ Acknowledgments
The author would like to thank Maggie Miller, Imogen Montague, Mark Powell, Anthony Conway, Evan Scott, Aru Mukherjea, and Seungwon Kim for their generous help and advice. Special thanks to Malcolm Gabbard for suggesting this question, to Mark Powell for his helpful comments on an earlier draft of this paper, and to Imogen Montague, for pointing out and subsequently helping to fix a mistake in an earlier draft.
The author received partial support from the Simons Foundation collaboration grant on New Structures in Low-dimensional topology while writing this paper.
1. Preliminaries on annuli and periodic 2-knots
1.1. Periodic 2-knots
@@
An action $\rho$ of $\mathbb{Z}/d\mathbb{Z}$ on the standard $m$-ball $B^{m}$ is linear if it is conjugate to the action of an $m$-dimensional orthogonal representation of $\mathbb{Z}/d\mathbb{Z}$ on $B^{m}$. An action of $\mathbb{Z}/d\mathbb{Z}$ on $S^{m-1}$ is linear if it is conjugate to the action of such a representation on $B^{m}$ restricted to the boundary.
The actions we consider in this paper are the the linear rotation action $\sigma$ of $\mathbb{Z}/d\mathbb{Z}$ on $S^{2}$ fixing two points, and the linear rotation action $\rho$ of $\mathbb{Z}/d\mathbb{Z}$ on $B^{5}$ fixing an unknotted 3-ball such that its restriction to the boundary $S^{4}$ fixes an unknotted 2-sphere. Restricting to linear actions simplifies the argument, although we hope to extend to nonlinear actions in future work.
Before proceeding, let us establish some notation for manifolds with cyclic group actions. We will frequently need to pass between equivariant manifolds and their quotients by $\rho$. Throughout this paper, we denote manifolds which carry an action by $\rho$ with tildes, to distinguish them from their quotients by $\rho$; e.g. $\tilde{Y}\subset\tilde{B}^{5}$ becomes $Y\subset B^{5}$ after taking the quotient by $\rho$. (Because $\rho$ is always a linear cyclic action by rotation, the quotient of $\tilde{B}^{5}$ by $\rho$ is always diffeomorphic to $B^{5}$, and the quotient of $\tilde{S}^{4}$ by $\rho$ is always diffeomorphic to $S^{4}$.) To recover the original manifold, we may take the $d$-fold cyclic branched cover over the image of the fixed set under the covering map. For this reason, we refer to the image of $\operatorname{Fix}(\rho)$ under the covering map as the branch set and denote it $\operatorname{Branch}(\rho)$. Objects contained in the fixed set of $\rho$ will be denoted with hats, to distinguish them from their images under the quotient map (e.g. $\tilde{Y}\cap\operatorname{Fix}(\rho)=\hat{A}$, but in the quotient $Y\cap\operatorname{Branch}(\rho)=A$).
For convenience, we give a general definition for embeddings which are equivariant with respect to our desired type of action in all dimensions.
@@
[…]
$\psi\circ\sigma=\rho\circ\psi.$
In the case that $\psi(\tilde{M})\cap\operatorname{Fix}(\rho)$ is a single, smoothly embedded copy of $S^{n-2}$, or a single properly embedded copy of $D^{n-2}$ in the case that $\tilde{M}$ has boundary, we say that $\psi$ is simply $d$-periodic.
Some abuses of this notation: where it is not explicitly relevant, we will often drop the prefix $d$- and describe embeddings just as periodic or simply periodic. Furthermore, we will also often describe the image of periodic embeddings as periodic; i.e., when we say that a 2-knot $\tilde{S}\subset S^{4}$ is periodic, we mean that there is an embedding $\psi\colon S^{2}\to S^{4}$ whose image is $\tilde{S}$, and that $\psi$ is periodic for some period $d$.
Following this definition, we see that a 2-knot $\tilde{S}$ in $S^{4}$ is $d$-periodic if it satisfies
$\psi\circ\sigma=\rho\circ\psi,$
with $\operatorname{Fix}(\rho)$ an unknotted 2-sphere. In this case, since the restriction of $\rho$ to $\tilde{S}$ is orientation preserving, we see by the Lefschetz fixed point theorem that $S\cap\operatorname{Fix}(\rho)$ consists of exactly two points. Such spheres arise naturally in common constructions of 2-knots: all $d$-twist spun knots can be constructed to have this symmetry about their spin axis.
@@
All periodic 2-knots are simply periodic as noted above, but the same is not true of surface knots in $S^{4}$. For instance, the unknotted torus in $S^{4}$ admits two different order 2 periodic symmetries: one in which one of the $S^{1}$ factors of the torus forms a meridian for $\operatorname{Fix}(\rho)$, and another that restricts to the hyperelliptic involution. The former action is free, while the latter fixes four points.
Note that often, the intersection of a simply periodic manifold $\tilde{M}$ with the ambient fixed set can be knotted in $\tilde{M}$ (rel. boundary, if applicable). This will be of particular importance to us.
1.2. Montesinos annuli
As alluded to in the introduction and explained in detail in Section 5, we study the equivariant concordance of periodic 2-knots by studying the concordance of their quotients by $\rho$ relative to the fixed set of the action. This concordance question is then equivalent to that of relative concordance of certain annuli in $S^{1}\times B^{3}$. Because of this, the majority of the paper takes place in the setting of these annuli, which we dub Montesinos annuli.